英語 (English)

基盤英語 (Basic English) (工 ((機 A)1 年))

鎌田スザーン · 非常勤講師 1 単位前期月 7 ・ 8

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

【授業の目的】In today’s globalized society, the ability to communicate in has become increasingly important to gain knowledge of world events and the ability to process technical information. In this class, students will participate in a variety of guided activities meant to foster communication skills. The emphasis will be on communication, as opposed to grammar.

【授業の概要】This class will be conducted entirely in English. You will be expected to speak, listen, read and write individually, in pairs and in groups. All students will take part in a group project and give 3-minute speeches. There will be a final test at the end of the course.

【キーワード】 English communication 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】To become better at and more comfortable with using English for

communication. 【授業の計画】

1. Meeting and Greeting

2. What is it?

3. Describing Japanese things

4. Book discussion

5. Group Project

6. Group Project

7. Group Project

8. Directions

9. Old Wives’ Tales

10. Making plans

11. Find Someone Who

12. Three minute speeches

13. Three minute speeches

14. Game Day

15. Final test 16. Class Review 【教科書】 No textbook will be used. Students are expected to bring paper, writing

implements and a dictionary to each class. 【参考書等】 [参考資料] 【成績評価の方法】 Grading will be based on attendance and participation (25%), group

project ( 25%), three-minute speech (25%) and final test (25%). 【再試験の有無】有【受講者へのメッセージ】 I hope that students will learn to enjoy using English.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220844 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ 鎌田 . (オフィスアワー: suekamata@msn.com)

English

Basic English (工 ((機 A)1 年))

名 Kamada · Part-time Lecturer 1 unit 前期月 7 ・ 8

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

Target⟩ In today’s globalized society, the ability to communicate in has become increasingly important to gain knowledge of world events and the ability to process technical information. In this class, students will participate in a variety of guided activities meant to foster communication skills. The emphasis will be on communication, as opposed to grammar.

Outline⟩ This class will be conducted entirely in English. You will be expected to speak, listen, read and write individually, in pairs and in groups. All students will take part in a group project and give 3-minute speeches. There will be a final test at the end of the course.

�������⟩ English communication Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ To become better at and more comfortable with using English for

communication.

Schedule⟩

1. Meeting and Greeting

2. What is it?

3. Describing Japanese things

4. Book discussion

5. Group Project

6. Group Project

7. Group Project

8. Directions

9. Old Wives’ Tales

10. Making plans

11. Find Someone Who

12. Three minute speeches

13. Three minute speeches

14. Game Day

15. Final test 16. Class Review Textbook⟩ No textbook will be used. Students are expected to bring paper, writing

implements and a dictionary to each class. Reference⟩ [参考資料] Evaluation Criteria⟩ Grading will be based on attendance and participation (25%),

group project ( 25%), three-minute speech (25%) and final test (25%). Re-evaluation⟩ 有 Message⟩ I hope that students will learn to enjoy using English.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220844 Contact (Office-Hour, Room, E-mail)⟩

⇒ Kamada . (Office Hour: suekamata@msn.com)